Two satellites A and B go around a planet in circular orbits having radii 4R and R,respectively. If the speed of satellite A is 3V,the speed of satellite B will be

Grade:10

1 Answers

Arun
6 years ago
F=GMm/R2,
now force of Gravitation provide centripetal force
therefore mv2/R=GMm/R2
v=[GM/R]1/2

Now, according to question
3v=[GM/4R]1/2
3v=1/2[GM/R]1/2  ----------(1)

and
let the velocity of the other planet be x,
x=[GM/R]1/2 -------------(2)

from (1) and (2)

3v=x/2,
x=6v

​hence solved
 
Regards
Arun (askIITians forum expert)